# A Hybrid Algorithm for Systems of Non-interacting Particles

### v2

@@ 3.2 Adaptive hybrid in multiple dimensions
We note that essentially the same procedure can be used to initialize the problem. We start by defining the SPDE solution on the entire grid and using the error estimation to identify where the particle representation is needed. The only difference is that at initialization we initialize the particle region directly from a particle description instead of generating random particles based on the SPDE representation.
Although the basic outline of the time step is the same, there are a number of issues that arise because the particle region is now a union of rectilinear regions in multiple dimensions. As in the one-dimensional algorithm we first advance the SPDE over the entire domain, including cells covered by the particle domain. We note that, in practice, the SPDE region is decomposed into multiple rectilinear subregions. When implemented in this form, the boundary region of a given grid patch that is in the interior of the SPDE region is filled from the SPDE data on other grids that contain the boundary points. In addition, at faces at the intersection of two different patches, the random numbers used to compute the stochastic fluxes are synchronized so that the flux at the face is consistent between the two grids.
@@
The necessary boundary cells are filled based on data from the SPDE description. In particular, from the SPDE solution, we determine the number of particles in the region, using the same probabilistic approach discussed in one dimension. These particles are then assigned positions based on sampling from a uniform distribution over the cell, as indicated by the dark blue particles in Figure 1 for the two-dimensional case.
We note that when implemented to operate on rectilinear patches, some of the boundary cells for one patch () may correspond to particle cells in another patch (). In this case the data is copied from to so that the representations are consistent. Furthermore, some cells can be boundary cells for two different grids. In this case, we need to ensure that both the number of particles and their locations are the same on both patches. Finally, we must guarantee that both interior particle cells and boundary particle cells that appear in two different grid patches (including their boundary cells) use the same random number to advance the particles. (This is implemented by associated the random numbers as attributes of the particles so that realizations of the same particle in different grid patches are consistent.)
@@
is the number of particles that transferred from to . We can also define to be the net number of particles transferred from to in the particle update. The computation of is illustrated in Figure 1. Arrows that point from red particles inside the particle region to magenta particles in the boundary region generate a positive increment to for the boundary cell that contains the magenta particle. Arrows that point from dark blue particles to cyan particles inside the particle region decrement for the boundary cell containing the dark blue particle. We then define to be the set of particle cells that share a face with SPDE cell and is the set of particle cells that satisfy Eq. (13). We then correct the SPDE solution using
@@ 4 Computational examples
In this section, we present several computational examples relating to the behavior of the particle system. We will compare particle simulations, finite volume method, linearized Gaussian approximation and hybrid method. First, we examine the differences between the particle description and the SPDE description for small particles numbers. We note that, for the simple Euler-Muruyama discretization presented here, the temporal truncation error leads to an over-prediction of the variance that converges with first-order accuracy in . However, skewness and kurtosis are insensitive to the time step.
@@
As noted above, the SPDE description of the particle system is based on an assumption that there are enough particles locally in the system for the stochastic flux of particles to be approximated by white noise. As the particle number becomes small, this assumption breaks down. One clear difference is that the PDF of the number of particles per cell from a pure particle algorithm takes on discrete values, namely where is non-negative integer, while the PDF for the SPDE is continuous. However, there are additional differences. In Figure 2 we show numerical PDFs of systems with 500 and 100 particles at equilibrium in one-dimension discretized with 100 zones (), corresponding to an average of 5 and 1 particles per cell, respectively. For particle simulations we plot the PDF values as discrete points. For the SPDE we choose bins of size which correspond to the width of one particle per cell centered at discrete particle values and plot those values as a continuous distribution. The PDF from the particle simulations show asymmetry for the 500 particle case that becomes more significant with 100 particles. For both cases, the linearized Gaussian algorithm results in a Gaussian distribution that includes negative values, which becomes more significant. The finite volume algorithm captures some of the asymmetry of the particle algorithm, but the agreement is fairly poor for 100 particles. In addition, although not as severe as the linearized Gaussian algorithm, the finite volume algorithm also generates negative values for both 5 particles per cell and 1 particle per cell.
@@
To quantify the differences in these distributions, we examine higher-order moments of the distribution. In Figure 3, we compute skewness and kurtosis of the equilibrium distributions for each method as a function of the number of particles per cell, again using a one dimensional problem with 100 zones. With 50 particles per zone, all of the methods predict a skewness close to zero and a kurtsosis of about three, corresponding to a Gaussian distribution. The Gaussian algorithm retains the Gaussian character (by construction) for all values of particles per cell. For the particle algorithm and the finite volume method, deviations from Gaussian begin to appear at 20 particles per cell and become significant at 10 particles per cell. The particle algorithm and the finite volume algorithm agree reasonably well at 10 particles per cell but begin to show discrepancies at 5 particles per cell that become more pronounced at 1 particle per cell.
@@
These higher-order statistics show that accuracy of finite volume SPDE description decreases as the number of particles per zone decreases. This allows us to establish a quantitative relationship between the number of particles per cell and breakdown in the accuracy of the finite volume discretization. The data in Figure 3 suggest that finite volume discretization begins to show significant errors with fewer than 5-10 particles per zone, suggesting 10 particles per zone is a reasonable value to use as a “refinement” criterion.
@@
As an initial test of the hybrid, we consider a dynamic case in which particles are diffusing into an initially void region. We again consider a one-dimensional period case with 100 zones with for and elsewhere, corresponding to 20 particles per cell. For this test we fix the particle region to be one-cell wider in each direction than the region where there are initially no particles. In Figure 4 we present statistics for an ensemble of 20,000 simulations. The means and variances are in good agreement for all 4 methods. For the linearized Gaussian method, however, the skewness and kurtosis, which reflect a Gaussian distribution, are significantly different than the values for the particle algorithm, particularly in the low particle density region. The finite volume algorithm does somewhat better, however, the errors in the low particle density region are still significant. The hybrid algorithm captures the peaks in skewness and kurtosis at low particle densities, demonstrating that the hybrid algorithm can accurately model systems with regions of low particle density.
@@
Next we consider examples in multiple dimension that demonstrate the dynamic adaptation of the methodology. The first example is a two-dimensional case with initial conditions given by an ellipsoidal region with 15 particles per cell surrounded by a larger elliptical region without any particles embedded in a background with 30 particles per cell, as illustrated in Figure 5. The simulation is performed on a grid. For the hybrid simulation we tag cells with fewer than 5 particles per cell to be in the particle region. We compare the hybrid using the finite volume discretization with the finite volume discretization in Figures 5 and 6. The inner ellipse is initially outside the particle region but as the system evolves the density decreases and the particle region expands to include the entire inner ellipse. At the outer edge of the larger ellipse, particles diffuse into the region and the particle region shrinks. Thus, this example shows regions entering and leaving the particle region. The hybrid algorithm and the finite volume algorithm produce similar large-scale features; however, the finite volume algorithm generates negative values (indicated by white dots in the figure) whereas the solution from the hybrid algorithm remains positive.
@@
The final example is a three-dimensional case with initial conditions given by a spherical region with 15 particles per cell surrounded by a larger spherical region without any particles embedded in a background with 30 particles per cell, as illustrated in Figure 7. The simulation is performed on a grid. As in the two dimensional case, we tag cells with fewer than 5 particles per cell to be in the particle region for the hybrid algorithm. Comparison of the hybrid with the finite volume discretization is presented in Figures 7 and 8. The inner sphere is initially outside the particle region. As the system evolves the density decreases and the particle region begins to expand. At only a small region of the inner sphere is still be modeled with the finite volume method. By the particle region has grown to include the entire inner sphere. At the outer edge of the larger sphere, particles diffuse into the region and the particle region shrinks. The hybrid algorithm and the finite volume algorithm again produce similar large-scale features; however, the finite volume algorithm generates negative values (indicated by white dots in the figure) whereas the solution from the hybrid algorithm remains positive.
@@ 5 Conclusions and further work
In this work we have presented a hybrid method for simulating the dynamics of non-interacting particles. This approach combines the simulation of the Dean-Kawasaki equation with truncated noise and particle dynamics in low-density regions. In particular, we developed the refinement criteria by comparing higher-order statistics. Compared to prior work on hybrid algorithms for particle systems, such as [1, 2], we focused specifically on regimes where particle density becomes low. In these regimes, our numerical experiments have shown that incorporating a particle description of the system is crucial for accurately preserving higher-order statistics and positivity of the solution. Additionally, we have extended the method to higher spatial dimensions with dynamic adaptation.
A next step is to consider higher-order time discretization methods, such as the Milstein scheme or the BDF2 method. Note that BDF2 is expected to provide higher accuracy for SPDEs with small noise, as discussed in [6]. Since the noise factor in Eq. (3) is , we expect to achieve higher-order accuracy when the number of particles is sufficiently large. It would be interesting to compare this threshold with the number of particles per cell required to ensure correct higher-order statistics and preserve the positivity of the solution.
We also aim to prove error bounds for the FVM applied to the modified equation from [12] and obtain a rate of convergence for the time discretization. Given that the modified equation has a well-behaved solution, we can directly analyze the error between the moments of the discretized and continuous solution, rather than using a weaker norm. Our goal is to compare how the choice of parameters determined by the discretization error aligns with the parameters chosen in [12] for the error between the modified and Dean-Kawasaki equations, as well as with the parameters obtained from simulations using the hybrid method. This type of analysis would allow us to link error estimates for the hybrid to the refinement criteria.
As observed in the simulations, it is not just the negativity of the solution that determines the choice of the particle description, but also the difference of the higher order statistics. An interesting question in this regard is how would the development of full discretization method of the SPDE that preserves the positivity of the solution influence the refinement criteria that we observed using standard FVM and Euler-Maruyama scheme.
The computational framework for hybrid algorithms developed here can be extended to a broader range of problems. As a first step in that direction, we plan to extend this hybrid method to interacting particle systems, incorporating an additional term in (3), where is the interaction potential. We will explore different interaction potentials, such as Coulomb or Leonard-Jones type, and study how these choices affect the refinement criterion.
@@
Another promising application of the presented method is to study particles immersed within a surface, particularly for practical problems where geometry is crucial, as discussed in [22]. Our goal is to develop a hybrid method that incorporates the role of geometry by simulating the dynamics of Brownian particles on a surface, solving the corresponding stochastic equation on that surface, and synchronizing these processes while accounting for the geometric factors.

### v3

@@ 3.2 Adaptive hybrid in multiple dimensions
We note that essentially the same procedure can be used to initialize the problem. We start by defining the SPDE solution on the entire grid and using the error estimation criterion to identify where the particle representation is needed. The only difference is that at initialization we have the option to initialize the particle region directly from a particle description instead of generating random particles based on the SPDE representation.
Although the basic outline of the time step is the same, there are a number of issues that arise because the particle region is now a union of rectilinear regions in multiple dimensions. As in the one-dimensional algorithm we first advance the SPDE over the entire domain, including cells covered by the particle domain. We note that, in practice, the SPDE region is decomposed into multiple rectilinear subregions. When implemented in this form, the boundary region for a given grid patch that is in the interior of the SPDE region is filled from the SPDE data on other grids that contain the boundary points. In addition, at faces at the intersection of two different patches, the random numbers used to compute the stochastic fluxes are synchronized so that the flux at the face is consistent between the two grids.
@@
The necessary boundary cells are filled based on data from the SPDE description. In particular, from the SPDE solution, we determine the number of particles in the region, using the same probabilistic approach discussed in one dimension. These particles are then assigned positions based on sampling from a uniform distribution over the cell, as indicated by the dark blue particles in Figure 2 for the two-dimensional case. As in one dimension, the sampling need to be modified when there is an exernal potential.
We note that when implemented to operate on rectilinear patches, some of the boundary cells for one patch () may correspond to particle cells in another patch (). In this case the data is copied from to so that the representations are consistent. Furthermore, some cells can be boundary cells for two different grids. In this case, we need to ensure that both the number of particles and their locations are the same on both patches. Finally, we must guarantee that both interior particle cells and boundary particle cells that appear in two different grid patches (including their boundary cells) use the same random numbers to advance the particles. (This is implemented by associating the random numbers as attributes of the particles so that realizations of the same particle in different grid patches are consistent.)
@@
is the number of particles that transferred from to . We can also define to be the net number of particles transferred from to in the particle update. The computation of is illustrated in Figure 2. Arrows that point from dark red squares inside the particle region to light red diamonds particles in the boundary region generate a positive increment to for the boundary cell that contains the light red diamonds particle. Arrows that point from dark blue triangles particles to light blue circles inside the particle region decrement for the boundary cell containing the dark blue triangle particle. We then define to be the set of particle cells that share a face with SPDE cell and is the set of particle cells that satisfy Eq. (14). We then correct the SPDE solution using
@@ 4 Computational examples
In this section, we present several computational examples relating to the behavior of the particle system. We will compare particle simulations with the finite volume method, the linearized Gaussian approximation and the hybrid method. The initial tests, which consider problems without an external potential, are designed to characterize the properties of the methodology in a simple setting. These initial simulations are performed on a periodic domain. The last example incorporates an external potential to illustrate the application of the methodology to rare event simulation.
@@
First, we examine the differences between the particle description and the SPDE description for small particles numbers. We note that, for the simple Euler-Maruyama discretization presented here, the temporal truncation error leads to an over-prediction of the variance that converges with first-order accuracy in . However, skewness and kurtosis are insensitive to the time step.
@@
As noted above, the SPDE description of the particle system is based on an assumption that there are enough particles locally in the system for the stochastic flux of particles to be approximated by white noise. As the particle number becomes small, this assumption breaks down. One clear difference is that the PDF of the number of particles per cell from a pure particle algorithm takes on discrete values, namely where is non-negative integer, while the PDF for the SPDE is continuous. However, there are additional differences. In Figure 3 we show numerical PDFs of systems with 500 and 100 particles at equilibrium in one-dimension discretized with 100 cells () corresponding to an average of 5 and 1 particles per cell, respectively, using a time step , which is small enough to mitigate the over-prediction of variance discussion above. For particle simulations we plot the PDF values as discrete points. For the SPDE we choose bins of size which correspond to the width of one particle per cell centered at discrete particle values and plot those values as a continuous distribution. The PDF from the particle simulations show asymmetry for the 500 particle case that becomes more significant with 100 particles. For both cases, the linearized Gaussian algorithm results in a Gaussian distribution that includes negative values, which becomes more significant as the number of particles decreases. The finite volume algorithm captures some of the asymmetry of the particle algorithm, but the agreement is fairly poor for 100 particles. In addition, although not as severe as the linearized Gaussian algorithm, the finite volume algorithm also generates negative values for both 5 particles per cell and 1 particle per cell.
@@
To quantify the differences in these distributions, we examine higher-order moments of the distribution. In Figure 4, we compute skewness and kurtosis of the equilibrium distributions for each method as a function of the number of particles per cell, again using a one dimensional problem with 100 cells. With 50 particles per cell, all of the methods predict a skewness close to zero and a kurtosis of about three, corresponding to a Gaussian distribution. The Gaussian algorithm retains the Gaussian character (by construction) for all values of particles per cell. For the particle algorithm and the finite volume method, deviations from Gaussian begin to appear at 20 particles per cell and become significant at 10 particles per cell. These deviations reflect the underlying multinomial distribution for a finite size domain. For comparison purposes in Figure 4 we include the skewness and kurtosis for the Poisson distribution, which approximates the multinomial distribution for large numbers of cells. The particle algorithm and the finite volume algorithm agree reasonably well at 10 particles per cell but begin to show minor discrepancies at 5 particles per cell that become more pronounced at 1 particle per cell.
@@
These higher-order statistics show that accuracy of finite volume SPDE description decreases as the number of particles per cell decreases. This allows us to establish a quantitative relationship between the number of particles per cell and breakdown in the accuracy of the finite volume discretization. The data in Figure 4 suggest that finite volume discretization begins to show significant errors with fewer than 5-10 particles per cell, suggesting 5-10 particles per cell is a reasonable value to use as a “refinement” criterion. We note that this estimate is somewhat conservative since the criterion is based on the average number of cells rather than the local variability. Thus, we expect that the threshold can be set toward the lower end of the range.
Refinement criteria based on higher-order statistics is different than what is typically used in hybrid algorithms for fluids such as those based on DSMC / continuum. In those cases, criteria based on gradients are typically used, see, for example, [37]. The use of gradients in the fluid hybrid setting is based on breakdown of the continuum model [7, 21]. However, tests on the particle system suggest that the SPDE description remains accurate in the presence of strong gradients provided that there are a sufficient number of particles per cell in the region of the gradient.
@@
As an initial test of the hybrid, we consider a dynamic case in which particles are diffusing into an initially void region. We again consider a one-dimensional periodic case with 100 cells with for and elsewhere, corresponding to 20 particles per cell. For this test we fix the particle region to be one-cell wider in each direction than the region where there are initially no particles. In Figure 5 we present statistics for an ensemble of 20,000 simulations. The means and variances are in good agreement for all 4 methods. For the linearized Gaussian method, however, the skewness and kurtosis, which reflect a Gaussian distribution, are significantly different than the values for the particle algorithm, particularly in the low particle density region. The finite volume algorithm does somewhat better, however, the errors in the low particle density region are still significant. The hybrid algorithm captures the peaks in skewness and kurtosis at low particle densities, demonstrating that the hybrid algorithm can accurately model systems with regions of low particle density.
@@
Next we consider examples in multiple dimension that demonstrate both the dynamic adaptivity and the positivity preservation property of the methodology. The first example is a two-dimensional case with initial conditions given by an ellipsoidal region with 15 particles per cell surrounded by a larger elliptical region without any particles embedded in a background with 30 particles per cell, as illustrated in Figure 6. Simulations for the finite volume method and the hybrid algorithm were performed on a grid. (We note that at this resolution 15 particles per cell corresponds roughly to a number density of .) For the hybrid simulation we tag cells with fewer than 5 particles per cell to be in the particle region, which is near the low end of the range of criteria discussed earlier. We present time sequences for finite volume discretization and the hybrid in Figures 6 and 7, respectively. In the hybrid algorithm, the inner ellipse is initially outside the particle region but as the system evolves the density decreases and the particle region expands to include the entire inner ellipse. At the outer edge of the larger ellipse, particles diffuse into the region and the particle region shrinks. Thus, this example illustrates the dynamic adaptivity of the hybrid with differ parts of the domain entering and leaving the particles region during the simulation. The hybrid algorithm and the finite volume algorithm produce similar large-scale features; however, the finite volume algorithm generates negative values (indicated by white dots in the figure) whereas the solution from the hybrid algorithm remains positive.
@@
Our next example is a three-dimensional case with initial conditions given by a spherical region with 15 particles per cell surrounded by a larger spherical region without any particles embedded in a background with 30 particles per cell, as illustrated in Figure 8. Simulations with the finite volume method and the hybrid algorithn were performed on a grid. At this resolution 15 particles per cell corresponds to a number density of roughly . As in the two dimensional case, for the hybrid we again tag cells with fewer than 5 particles per cell to be in the particle region. Results for the finite volume discretization and the hybrid are presented in Figures 8 and 9. The inner sphere is initially outside the particle region. As the system evolves the density decreases and the particle region begins to expand. At only a small region of the inner sphere is still be modeled with the finite volume method. By the particle region has grown to include the entire inner sphere. At the outer edge of the larger sphere, particles diffuse into the region and the particle region shrinks. The hybrid algorithm and the finite volume algorithm again produce similar large-scale features; however, the finite volume algorithm generates negative values (indicated by white dots in the figure) whereas the solution from the hybrid algorithm remains positive.
Our final numerical example illustrates the application of the methodology to a problem with an external potential. The modifications required to incorporate an external potential into the algorithm are discussed in the Appendix A. Here we consider a potential of the form
which has minima at and . When in Eq. (1) is sufficiently small, the transition from to becomes a rare event. Problems of this type arise in the calculation of rate coefficients in physics, chemistry and biology, originating with early work of Kramers [31]. Schuss [34] presents a diffusion model for chemical reactions, where the confining potential well consists of a series of holes and barriers and the particle dynamics is given by the Langevin dynamics of the type (1). Hänggi et al. [25] discusses a number of developments in this area. A seminal paper of Jordan, Kinderlehrer and Otto [28] investigates Fokker–Planck equations where the drift term is given by the gradient of a potential, which corresponds to the dynamics described by (1). Similar types of problems also arise in large-deviation theory, see Berglund [6] and are directly relevant to simulations in molecular dynamics [35].
We set and and consider the case where . We initialize the simulation with roughly particles in the subregion distributed based on the equilibrium distribution
restricted to the subregion ( is the normalization constant). We discretize the system on a grid with homogeneous Dirichlet boundary conditions and simulate for 100,000 time steps with using the finite volume scheme, particles dynamics and the hybrid method. See Appendix B for the treatment of boundary conditions. Results from the three methods are shown in Figure 10. The presence of a potential leads the bulk of the particles being localized to small regions of the domain with very low particle density in the remainder of the domain. This results in a loss of positivity for the finite volume scheme in a significant fraction of the domain as indicated in Figure 10.
In Figure 11, we reduce the range on the images to highlight the particles that have crossed the energy barrier. We consider particles whose location is greater than 0.6 to be in the potential well at . Here 0.6 corresponds to the location where the potential is half the barrier height. We measured the number of particles that had crossed the barrier into the potential well at for a small ensemble of ten simulations with each method. For the particle algorithm and the hybrid method we obtained means at the final time of 388 and 384 particles respectively, which are within statistical error bars for the small ensemble. The finite volume algorithm gave a mean of 370 particles, which is a statistically significant under prediction. A much larger ensemble from a one-dimensional case showed similar results with the finite volume method under predicting the number of particles crossing the barrier while hybrid and particle simulations predict means that differ by less than 1%.
As noted above, when we generate particles from the SPDE representation during regridding or when filling ghost cells, we do not assume that the particles are uniformly distributed in the cell when there is an external potential. Instead we sample grid locations from a local approximation to the equilibrium distribution. To illustrate the importance of using the appropriate measure for this operation, we performed a small ensemble of simulations with the external potential in which we defined particle locations for regridding and filling ghost cells using a uniform distribution. For this ensemble, the mean number of particles that crossed to the potential well at was 555, representing a dramatic over-estimation of the number of particles crossing the potential barrier.
@@ 5 Conclusions and further work
In this work we have presented a hybrid method for simulating the dynamics of non-interacting particles with the example that also includes external potential. This approach combines the simulation of the Dean-Kawasaki equation with truncated noise and particle dynamics in low-density regions. In particular, we developed the refinement criteria by comparing higher-order statistics. Compared to prior work on hybrid algorithms for particle systems, such as [1, 2], we focused specifically on regimes where particle density becomes low. In these regimes, our numerical experiments have shown that incorporating a particle description of the system is crucial for accurately preserving higher-order statistics and positivity of the solution. Additionally, we have extended the method to higher spatial dimensions with dynamic adaptation and we have provided a numerical example that includes an external potential.
An important future direction for this work is the extension of the hybrid framework to include interactions between the particles. There has been a substantive amount of work on SPDE models that include interactions between the particles, see, for example, [36, 26, 27, 8]; however, there has been essentially no work on development of hybrid algorithms of the type considered here. Developing hybrids that include particle interactions would enable a broad range of applications in physics, chemistry, biology and the social sciences. Unfortunately, incorporating these types of models into a hybrid poses a number of challenges. First, one needs to be able to compute the interaction between parts of the solution represented by particles and parts of the solution represented by number density. The second major issue is how to generate particles consistent with the SPDE state for regriding and to fill ghost cells. As shown here for an external potential, the numerical results for the hybrid are extremely sensitive to how particles are generative. The solution to either of these problems is not known.
@@
Another promising application of the methodology would be to study particles immersed within a surface for problems where geometry is crucial, as discussed in [33]. Here the goal would be to develop a hybrid method that incorporates the geometry by simulating the dynamics of Brownian particles on a surface, solving the corresponding stochastic equation on that surface, and synchronizing these processes while accounting for the geometric factors.
There are also a number of more mathematical research topics suggested by this work. One area would be to prove error bounds for the FVM applied to the modified equation from [15] and obtain a rate of convergence. Given that the modified equation has a well-behaved solution, one could directly analyze the error between the moments of the discretized and continuous solution, rather than using a weaker norm. The goal would be to compare how the choice of parameters determined by the discretization error aligns with the parameters chosen in [15] for the error between the modified and Dean-Kawasaki equations, as well as with the parameters obtained from simulations using the hybrid method. This type of analysis would allow us to link error estimates for the hybrid to the refinement criteria.
There are also opportunities to improve the basic SPDE methodology. Of particular interest would be higher-order time discretization methods, such as the Milstein scheme or the BDF2 method. Note that BDF2 is expected to provide higher accuracy for SPDEs with small noise, as discussed in [9]. Since the noise factor in Eq. (4) is , we expect to achieve higher-order accuracy when the number of particles is sufficiently large. It would be interesting to compare this threshold with the number of particles per cell required to ensure correct higher-order statistics and investigate the whether improvements to the temporal integrator impacts the refinement criterion.
Appendix A: Discretization including an external potential
In this appendix we briefly discuss the modifications needed to include an external potential into the dynamics. We will retain the simple Euler-Maruyama temporal discretization that was used in the absence of an external potential. The discretization of the particle motion then becomes
where ’s are vectors of independent sample from a normal distribution.
For the SPDE, we add an additional flux corresponding to the potential term . Specifically, we define
We then define the total flux to be
With these modifications, the hybrid algorithm is essentially unchanged. The only modification that is needed is in the algorithm for defining a particle presentation in a cell that is consistent with the SPDE solution. When there is an external potential, the equilibrium distribution of particles is no longer uniform. Instead it is given by
where is the position in space and is a normalization factor. (For the conditions considered here, is large enough at the domain boundary that the support of is essentially contained inside the domain.) When we need to define particles in a given cell consistent with the SPDE solution in that cell, we want to conditionally sample form Eq. (15) instead of assuming the particles are uniformly distributed. This can be important when the potential changes significantly within a given cell. Here, we have constructed a local approximation to equilibrium potential based on the potential gradient at the cell center. This approximation allows us to account for variability within the cell but is straightforward to sample.
For both particle update and SPDE integration, one needs to consider whether that addition of the external potential necessitates a smaller time step. In particular, we still need to constrain a particle to not move too far during a time step and ensure the time is small enough that the SPDE discretization is stable and accurate. For the case considered here we have not found reduction in time step necessary.
Appendix B: Discretization of boundary conditions
For Dirichlet boundary conditions, for the SPDE we impose the boundary condition on the face that coincides with the boundary. The diffusive component of the flux is then given by
where is the Dirichlet value on the boundary at . This formula reflects the reduction in the distance between the points to . The stochastic flux at the boundary is then multiply by to preserve fluctuation balance at the boundary. The potential flux is simply evaluated using the boundary condition.
For the particle regions that touch the boundary, the boundary condition is represented by a reservoir of particles outside the domain, as done in [1]. The number of particles in the reservoir is obtained by sampling from a Poisson distribution with mean given by the boundary condition. The location of the particles in the reservoir are sampled from the appropriate equilibrium distribution. After the particles are advanced, particles outside the domain are discarded.